OpenStudy (anonymous):

what line does the point (7, 1) lie?

OpenStudy (anonymous):

are you given a choice of lines on which it might lie? there are an infinite amount of lines through any point in the plane

OpenStudy (anonymous):

y=mx - ym + 1 x-intercept = 7-1/m y-intercept = 1-7m

OpenStudy (anonymous):

yes i am they are Y=5X +4 Y= X -10 Y=-4X + 3 AND Y=-X +8

OpenStudy (anonymous):

Then, -5x+y-4=0

OpenStudy (anonymous):

whew now we have a question \((x,y)\) in your case is \((7,1)\) replace \(x\) by \(7\) and \(y\) by \(1\) and see if the equation is true or false

OpenStudy (anonymous):

for example, lets check \[y=5x +4\] \[1=5\times 7+4\] no way lets check another

OpenStudy (anonymous):

\[Y= X -10\] \[1=7-10\] nope

OpenStudy (anonymous):

\[Y=-4X + 3\] think it will work here?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

is it clear what i am doing or not?

OpenStudy (anonymous):

NO WAY

OpenStudy (anonymous):

SORRY

OpenStudy (anonymous):

ok lets go slow you want to know which of these four lines contains the point \[(\color{red}{7},\color{blue}{1})\]

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

if the point is on the line, that means when you replace \(\color{red}{x}\) by \(\color{red}{7}\) and \(\color{blue}{y}\) by \(\color{blue}{1}\) the equation should be true, but which i mean the number on the left (which is a \(\color{blue}{y}\) and therefore is \(\color{blue}{1}\) should be equal to the number on the right

OpenStudy (anonymous):

Can you please just give me the answer?

OpenStudy (anonymous):

Pretty please with sprinkles on top

OpenStudy (anonymous):

so your job is to replace \(\color{red}{x}\) by \(\color{red}{7}\) and \(\color{blue}{y}\) by \(\color{blue}{1}\) and see what you get

OpenStudy (anonymous):

sure lets check the last one

OpenStudy (anonymous):

what do i do when i replace them

OpenStudy (anonymous):

compute the number you get on the right, and see if it is equal to 1

OpenStudy (anonymous):

can u just give me the answer

OpenStudy (anonymous):

\(\color{blue}{1}=-\color{red}7 +8\)

OpenStudy (anonymous):

true or false?

OpenStudy (anonymous):

ok now what one of the choices i previously typed is it 1 2 3 or the fourth one

OpenStudy (anonymous):

\(\color{blue}Y=-\color{red}X +8\)

OpenStudy (anonymous):

it's either true or false:)

OpenStudy (anonymous):